Find k if 3x+ky=1 is perpendicular to y+2x=-5
Answer by lwsshak3(11628)

Find k if 3x+ky=1 is perpendicular to y+2x=-5
**
y+2x=-5
y=-2x-5
slope=-2
..
3x+ky=1
ky=-3x+1
y=-3x/k+1/k
slope=-3/k=1/2 (negative reciprocal of slope of first equation)
k=-6